The government is considering levying a tax of $50 per unit on suppliers of either concert tickets or bus passes. The supply curve for each of these two goods is identical, as you can see on each of the following graphs. The demand for concert tickets is shown by  The government is considering levying a tax of $50  (on the first graph), and the demand for bus passes is shown by    (on the second graph).
Answer #1

If the government taxes concert tickets, it receives $3000($12,000,$9,000,$3,000or $6,000) in the revenues and causes $7500($7,500,$3,000,$1,250or $1,000) of deadweight loss. If taxes bus passes, it receives $12000($9,000,$15,000,$12,000or$6,000) in tax reveues and causes $3000($7,500,$1,000,$1,250 or$3,000) of dead weight loss. If the government wants to get more tax revenue with less deadweight loss, it should taxBUS PASSES(bus passes or concert tickets). This is because, all else being equal, taxes on a good with relatively lower(lower or higher) price elasticity of demand generate more tax revenue and less deadweight loss.

PRICE (Dollars per ticket] 120 Govt Revenue + Tax 100 Deadweight Loss 80 60 D1 40 20 00 200 300 400 500 600 QUANTITY (Ticket

in the 1st case the tax revenue is area of the green rectangle=(70-60)x50=$3000

dead weight loss is area of the black one=0.5x(70-60)x250+0.5x(60-10)x250=$7500

PRICE (Dollars per pass 120 Govt Revenue S+ Tax 100 Deadweight Loss 80 60 40 20 100 200 300 400 500 600 QUANTITY (Passes Hel

in the 2nd case tax revenue is area of green rectangle=(100-40)x200=$12000

and dead weight loss is =0.5x100x(100-60)+0.5x100x(60-40)=$3000